• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Saha

    Log entry time 16:00:49 on March 9,2001

    Entry number 58158

    keyword=Shift Summary

    Shift Summary
    0800: At Sing12 setting. BeO target
    Checking Beam spot and harp scan. Harp scan screen needs to be rebooted
    0845: Harp scan 1248 and BeO check. Beam spot x=0, y=2mm (up)
    0850: Moving to LD2 target. 30 uA cw with raster on. FFB on
    0855-0910: Run 2664 Sing 12 setting. LD2 target (Rad. out), 0.9M events
    0912-0927: Run 2665 Sing 12 setting. LD2 target (Rad. out), 0.9M events
    0935-0949: Run 2666 Sing 12 setting. LD2 target (Rad. out), 0.9M events
    Radiator in 6%
    0953-0957: Run 2667 Sing 12 setting. LD2 target (Rad. 6%), 0.3M events
    Deadtime high ~36%. Lowering current to 20 uA. Change prescale.
    0958-1009: Run 2668 Sing 12 setting. LD2 target (Rad. 6%), 0.9M events
    Deadtime reduced to 26%
    1010-1020: Run 2669 Sing 12 setting. LD2 target (Rad. 6%), 0.9M events
    1020-1032: Run 2670 Sing 12 setting. LD2 target (Rad. 6%), 0.9M events
    1033-1044: Run 2671 Sing 12 setting. LD2 target (Rad. 6%), 0.9M events
    1044-1055: Run 2672 Sing 12 setting. LD2 target (Rad. 6%), 0.9M events
    1057-1107: Run 2673 Sing 12 setting. LD2 target (Rad. 6%), 1 M events
    1108-1119: Run 2674 Sing 12 setting. LD2 target (Rad. 6%), 1 M events
    1119-1130: Run 2675 Sing 12 setting. LD2 target (Rad. 6%), 1 M events
    1132-1144: Run 2676 Sing 12 setting. LD2 target (Rad. 6%), 1 M events
    1145-1156: Run 2677 Sing 12 setting. LD2 target (Rad. 6%), 1 M events
    1157-1207: Run 2678 Sing 12 setting. LD2 target (Rad. 6%), 1 M events
    1208-1220: Run 2679 Sing 12 setting. LD2 target (Rad. 6%), 1 M events
    1221-1233: Run 2679 Sing 12 setting. LD2 target (Rad. 6%), 1 M events
    1234-1244: Run 2679 Sing 12 setting. LD2 target (Rad. 6%), 1 M events
    1245-1256: Run 2679 Sing 12 setting. LD2 target (Rad. 6%), 1 M events
    1256-1300: Run 2679 Sing 12 setting. LD2 target (Rad. 6%), 0.4 M events
    1300: Changing to new beam energy E=1.875 GeV
    Setting spectrometers to COIN20 settings
    Left arm: p = 1.3698 GeV/c, theta = 36.719 deg
    Right arm: p = -1.099 GeV/c, theta = 48.196 deg
    1500: Right arm Dipole stuck. Had to reboot right arm magnet IOC.